# VoxelKeypointFusion: Generalizable Multi-View Multi-Person Pose Estimation

### v1

@@ 1 Introduction
There exist several methods to estimate poses from persons, for example by attaching markers to the person’s body which can be tracked by special cameras, or by using normal cameras and estimating the pose from their images. While the marker-based approach is generally more accurate, it is also more complicated because the persons normally need to wear special clothes. This is not convenient for the persons and not always possible, for example, if one likes to track persons in public places, like a shopping store, or if special clothing is required, like in operation rooms. The marker-less approach, on the other hand, is much more convenient, but also more difficult to calculate, since the poses need to be estimated from the images only. Usually, multiple cameras are used to capture the scene from different viewpoints, since this makes the estimation more robust to occlusions and can provide more accurate results.
@@
Currently, almost all approaches only predict the main body joints, like eyes and nose or elbow and wrists, but for some applications this is not enough. For example, if one likes to analyze the actions of a worker in a human-robot collaboration, the movements of the fingers contain much more information than the movements of the wrists alone. Therefore, it is shown that the presented algorithm can not only generalize well to new datasets, but also to different keypoints, which allows the estimation of whole-body poses (see Figure 1 for example).
Figure 1: Example of multi-person whole-body pose estimation from multiple camera views, on the left the per-image 2D estimations, on the right their fused 3D poses.
The contributions of this work can be summarized as follows: (a) the first broad evaluation of the generalization capabilities of multi-view multi-person pose estimators across multiple unseen datasets, (b) the introduction of a new well generalizing algorithm, (c) the implementation of the first multi-view multi-person whole-body pose estimator (d) the introduction of a new library to simplify the handing of dataset related tasks.
@@ 2 Related Work I
In some cases, it might be possible to reduce the task of estimating human poses to the prediction of a single person only [10, 31, 6], which would greatly simplify the problem, because there is no need to match multiple joints to different persons having multiple possible options. However, since it allows a broader range of applications, the focus of this work is put on the multi-person algorithms. It is also assumed that the cameras have been calibrated in advance, a process that should be feasible in the majority of use cases.
@@ 3 VoxelKeypointFusion
Figure 2: Obtaining keypoints and person ids. First, the heatmaps of each view (a, schema for two cameras, the 3D-skeletons are drawn only for better visualization) are projected as beams into the voxelized room (b, five real cameras, only right hip joint visualized). Then peaks at overlapping beam positions are searched and retrieved as keypoint proposals. The peak proposals, here visualized as pink crosses, are projected into each image (c, with the two cameras from a, for better visualization they do not match to overlaps in b). The person association is then gathered from the person-id images (c, color-coded and color-paired only for better visualization, so every pixel with a red color has id=$1$). The person-ids of these points are extracted, in this case, one proposal has ids {$3$,$5$}, and the other two received no ids and thus are discarded. The same is done with the proposals from different joint types (so there could be a right shoulder with ids {$3$,$5$}). All proposals with the same id-set are then collected into a person group, which might be merged with other overlapping groups (for example if the left shoulder only got {$3$}).
In step (1) the keypoints of the 2D pose estimators are used to generate Gaussian heatmap images for each person in the view, which are then combined into a single heatmap. Basically one could also use heatmaps directly from the pose estimator, but not every model has them and it also would need extra implementations for every estimator. In step (2) images of the same shape are created, but instead of keypoint scores, they contain a person-id everywhere the heatmap score is not zero. The ids are generated by simply counting the total number of persons in all images, and every 2D person gets a different number. Those two steps can be seen on the two cube walls in Figure 2 (a+c). The voxel-based projection (default resolution $50mm$) in steps (3-5) is quite similar to [10, 25], except there is no learnable updating of the scores and only a single voxel-room is used (see Figure 2, b). The heatmap beams overlap at certain points, basically everywhere a joint is seen from at least two views, and sometimes with beams from other persons. The score of each voxel is calculated as the average of the heatmap beams, so overlapping beams lead to higher scores. Each voxel over a certain threshold (after non-maximum-neighbour-suppression) is selected as keypoint proposal. In step (6) each keypoint proposal is reprojected to the input views and obtains a set of corresponding person indices from the person-id images if they fall onto a 2D person keypoint (see Figure 2, c). If the proposal does not match any person, for example when beams from two different persons overlap somewhere in the space, it is discarded directly. The association to persons then happens in steps (7+8). In the perfect case, the keypoint proposals of one person all receive the same set of ids and can be clearly matched together. Since this is often not the case, due to occlusions by other persons, which lead to the assignment of wrong ids, groups where the majority of the person ids overlap are merged together, thus ignoring single wrong assignments. In steps (9-11) one person for each group is created by first taking the best proposal for each keypoint, to calculate an initial center, then improving the center by ignoring outliers far away, and then using this second center to drop all keypoints with too large distance in 3D space (above $1.3m$). Then the proposals for the torso joints with the best scores are taken to build an initial person. This person is extended with limbs, by first dropping far away outliers again (above $0.6m$ from the parent joint, since the limb lengths of a person are physically restricted) and assigning the remaining best keypoint proposal. At last any persons with too few keypoints (less than $3$) are dropped as incomplete, and the remaining persons are returned as the final result.
@@ 5 Dataset Generalization - Part I
Before training VoxelPose with synthetic labels and the new camera setup, a transfer with the original Panoptic [12] dataset was tried. Since this dataset has a large number of cameras, one training with five randomly chosen cameras in each training step, and another one with four fixed cameras with a similar room distribution as in Human36m, was executed, but both concepts did not improve the performance much. In difference to that, training with synthetically placed poses and their reprojections to the matching camera views greatly improved the performance. The synthetic training with Faster-VoxelPose significantly improved most metrics as well, especially the number of invalid predictions. A finetuning of MvP and TEMPO was not possible due to unpublished tools and missing documentation about how to generate their synthetic data inputs. SelfPose3d followed the concept of VoxelPose using data from the Panoptic dataset instead of self-supervised data for other datasets than Panoptic itself, but those parts were not published as well (the reported results in the paper were slightly worse compared to the ones of VoxelPose).
@@
The second evaluation was run with the Shelf & Campus [1] datasets. Shelf shows multiple persons assembling a shelf, from five different viewpoints, while Campus shows an outside scene with people walking and standing on a place in front of a building, watched by three cameras. Both datasets are commonly used for generalization tests and were also used in most models evaluated here. In difference to the other evaluations that normally only reported PCP, more metrics are used, and the PCP score is calculated as described above, and not by first averaging over the persons and then merging them together like usually, because this, on the one hand, results in a different weighting of errors since some persons occur more often than others, and on the other hand, allows a better comparison with other datasets.
In Table 3 it can be seen that the learnable approaches show similar performance on the Shelf dataset compared to the algorithmic ones. They are notably faster than most others, which mainly is caused by faster 2D pose estimators, but on the other hand, tend to predict more invalid persons. Note that in the dataset itself not all occurring persons are labeled, and therefore the percentage can not be zero, but the algorithmic models can detect all labeled persons while predicting a lower total number of persons. MvP and SelfPose3d show an especially poor performance in this regard, making them basically unusable for real-world applications. While VoxelPose profits from the synthetic training, the difference for Faster-VoxelPose is only marginal or even slightly worse in some metrics. The camera setup of Shelf is quite similar to the Panoptic dataset, which is a reason the directly transferred models show a decent performance. In all cases, the models were evaluated end-to-end, instead of using the provided external heatmaps created by unpublished tools, since the generalization of the complete model is the interesting part here.
@@ 6.2 VoxelKeypointFusion
For VoxelKeypointFusion a simple depth-based masking concept was implemented. After generating a point cloud from the depth images, all points that fall into one voxel are counted, and if there are more than a small threshold (to drop noisy points), this voxel is considered filled. The projection space is then masked with the filled depth voxels, to drop all projection voxels that do not match to a real object. An example can be seen in Figure 3. The depth views do not have to be aligned to the color views, but can be recorded from any direction and also come from a different number of sensors.
Figure 3: Example of depth masking with VoxelKeypointFusion. In the left image, the keypoint projection of four persons from five camera views into a voxelized room is shown, with one color for each joint type. This is the default input for the peak proposal calculation. On the right side, the projection was masked with the depth information, three of the persons in the center of the room are now clearly visible (the fourth is walking through the entrance), and no peaks can be proposed between the persons anymore. Parts of the sphere’s wall are still left, but since the keypoint projections do not overlap there, no peak proposals are generated from those.
@@ 8 Ablation studies
One limitation of VoxelKeypointFusion is, that since there is no neural refinement, the location accuracy is highly dependent on the voxel-resolution. As can be seen in Table 7, a lower voxel-resolution leads to better results, even though it does not directly scale (but $25mm$ is even better than the result with $50mm$ in Table 1). The voxel size also has a significant impact on the inference speed. As expected, with more cameras the results get better, but the inference time increases as well. With less cameras the performance drops, and the model predicts more invalid persons because overlaps where no persons stand are not pruned that much by different viewpoints that would see an empty space there. The number of persons, especially if there are many occlusions, has a notable impact on the inference speed as well. The total 3D time for Shelf is around $48ms$, for Campus around $56ms$, and for Panoptic, with fewer voxels but more total and occluded keypoints than Campus, around $100ms$.
@@ 10 Conclusion
In a direct comparison with VoxelPose, VoxelKeypointFusion works without any learnable part. Instead of the top-down approach of first detecting persons, then their joints, a bottom-up approach is used here, which first detects all joints, and then assembles them to persons. This is especially helpful in settings with stronger occlusion, because the risk of totally missing a person is lower. VoxelKeypointFusion also shows that all the necessary person information is already present in the 2D heatmaps, which currently is ignored by the learnable voxel-based approaches. One limitation is that since there is no neural refinement, the location accuracy is highly dependent on the voxel-resolution. In difference to previous line-based triangulation approaches, the voxel-based approach allows for multiple keypoints for each ray, which helps to detect occluded keypoints.

### v2

@@ 1 Introduction
There exist several methods to estimate poses from persons, for example by attaching markers to the person’s body which can be tracked by special cameras, or by using normal cameras and estimating the pose from their images. While the marker-based approach is generally more accurate, it is also more complicated because the persons normally need to wear special clothes. This is not convenient and also not always possible, for example, if one likes to track persons in public places, like a shopping store, or if special clothing is required, like in operation rooms. The marker-less approach, on the other hand, is much more convenient, but also more difficult to calculate, since the poses need to be estimated from the images only. Usually, multiple cameras are used to capture the scene from different viewpoints, since this makes the estimation more robust to occlusions and can provide more accurate results.
@@
Currently, almost all approaches only predict the main body joints, like eyes and nose or elbow and wrists, but for some applications, this is not enough. For example, if one likes to analyze the actions of a worker in a human-robot collaboration, the movements of the fingers contain much more information than the movements of the wrists alone. Therefore, it is shown that the presented algorithm can not only generalize well to new datasets, but also to different keypoints, which allows the estimation of whole-body poses (see Figure 1 for example).
Figure 1: Example of a multi-person whole-body pose estimation from multiple camera views (from the panoptic dataset [13]) with VoxelKeypointFusion, on the left the per-image 2D estimations, the right shows their fused 3D poses.
The contributions of this work can be summarized as follows: (a) the first broad evaluation of the generalization capabilities of multi-view multi-person pose estimators across multiple unseen datasets, (b) the introduction of a new well generalizing algorithm, (c) the implementation of the first multi-view multi-person whole-body pose estimator (d) the introduction of a new library to simplify the handling of dataset related tasks.
@@ 2 Related Work I
In some cases, it might be possible to reduce the task of estimating human poses to the prediction of a single person only [11, 32, 1, 7], which would greatly simplify the problem, because there is no need to match multiple joints to different persons having multiple possible options. However, since it allows a broader range of applications, the focus of this work is put on the multi-person algorithms. It is also assumed that the cameras have been calibrated in advance, a process that should be feasible in the majority of use cases.
@@ 3 VoxelKeypointFusion
Figure 2: Obtaining keypoints and person ids. First, the heatmaps of each view (a, schema for two cameras, the 3D-skeletons are drawn only for better visualization) are projected as beams into the voxelized room (b, five real cameras, only a single hip joint visualized). Then peaks at overlapping beam positions are searched and retrieved as keypoint proposals. The peak proposals, here visualized as pink crosses, are projected into each image (c, with the two cameras from a, for better visualization they do not match to overlaps in b or a). The person association is then gathered from the person-id images (c, color-coded and color-paired only for better visualization, so every pixel with a red color has id=$1$). The person-ids of these points are extracted, in this case, one proposal has ids {$3$,$5$}, and the other two received no ids and thus are discarded. The same is done with the proposals from different joint types (so there could be a right shoulder with ids {$3$,$5$}). All proposals with the same id-set are then collected into a person group, which might be merged with other overlapping groups (for example if the left shoulder only got {$3$}).
In step (1) the keypoints of the 2D pose estimators are used to generate Gaussian heatmap images for each person in the view, which are then combined into a single heatmap. Basically one could also use heatmaps directly from the pose estimator, but not every model has them and it also would need extra implementations for every estimator. In step (2) images of the same shape are created, but instead of keypoint scores, they contain a person-id everywhere the heatmap score is not zero. The ids are generated by simply counting the total number of persons in all images, and every 2D person is assigned a different number (the order does not matter). Those two steps can be seen on the two cube walls in Figure 2 (a+c). The idea is that the information of which joint belongs to which person is already present in the 2D pose estimations. The association of which person is which in the different views is not relevant here, joints of later 3D proposals only need not to be mixed up between different persons in the same view.
The voxel-based projection (default resolution $50mm$) in steps (3-5) is quite similar to [11, 26], except there is no learnable updating/refinement of the scores and only a single voxel-room is used (see Figure 2, b). The Gaussian heatmap beams overlap at certain points, basically everywhere a joint is seen from at least two views, and sometimes with beams from other persons. The score of each voxel is calculated as the average of the heatmap beams, so overlapping beams lead to higher scores, and well overlapping beams lead to higher scores than poorly overlapping ones. Each voxel over a certain threshold (after non-maximum-neighbor-suppression) is selected as keypoint proposal. The exact location is then adjusted with a weighted average of the neighboring voxels, to reach a sub-voxel precision.
In step (6) each keypoint proposal is reprojected to the input views and obtains a set of corresponding person indices from the person-id images if they fall onto a 2D person keypoint (see Figure 2, c). If the proposal does not match any person, for example when beams from two different persons overlap somewhere in the space, it is discarded directly. The association to persons then happens in steps (7+8). In the perfect case, the keypoint proposals of one person all receive the same set of ids and can be clearly matched together. Since this is often not the case, due to occlusions by other persons, which lead to the assignment of wrong ids, groups where the majority of the person ids overlap are merged together, thus ignoring single wrong assignments.
In steps (9-11) one person for each group is created by first taking the best proposal for each keypoint to calculate an initial center, then improving the center by ignoring outliers far away, and then using this second center to drop all keypoints with too large distance in 3D space (above $1.3m$). Then the proposals for the torso joints with the best scores are taken to build an initial person. This person is extended with limbs, by first dropping far away outliers again (above $0.6m$ from the parent joint, since the limb lengths of a person are physically restricted) and assigning the remaining best keypoint proposal. At last any persons with too few keypoints (less than $3$) are dropped as incomplete, and the remaining persons are returned as the final result.
In a direct comparison with VoxelPose, VoxelKeypointFusion works without any learnable part. Instead of the top-down approach of first detecting persons, then their joints, a bottom-up approach is used here, which first detects all joints, and then assembles them into persons. This is especially helpful in settings with stronger occlusion, because the risk of totally missing a person is lower. With the new concept of adding person-ids, already present information about the association of keypoints from the 2D detections is used, which allows skipping the person detection step. In difference to previous algebraic/line-based triangulation approaches, the voxel-based approach allows partially overlapping heatmap rays which helps to find a good center between slightly erroneous 2D predictions. The persons are also assembled directly in 3D space, which allows simpler outlier filtering steps.
@@ 5 Dataset Generalization - Part I
Before training VoxelPose with synthetic labels and the new camera setup, a transfer with the original Panoptic [13] dataset was tried. Since this dataset has a large number of cameras, one training with five randomly chosen cameras in each training step, and another one with four fixed cameras with a similar room distribution as in Human36m, was executed, but both concepts did not improve the performance much. In difference to that, training with synthetically placed poses and their reprojections to the matching camera views greatly improved the performance. The synthetic training with Faster-VoxelPose significantly improved most metrics as well, especially the number of invalid predictions. A finetuning of MvP and TEMPO was not possible due to unpublished tools and missing documentation about how to generate their synthetic data inputs. In their original paper TEMPO reported a transfer MPJPE of $63.0\,mm$ [5]. Reasons for the lower performance here can be differences in the data sampling and the published weights, which already showed a lower performance in the panoptic dataset. SelfPose3d followed the concept of VoxelPose using data from the Panoptic dataset instead of self-supervised data for other datasets than Panoptic itself, but those parts were not published as well (the reported results in the paper were slightly worse compared to the ones of VoxelPose).
@@
The second evaluation was run with the Shelf & Campus [2] datasets. Shelf shows multiple persons assembling a shelf, from five different viewpoints, while Campus shows an outside scene with people walking and standing on a place in front of a building, watched by three cameras. Both datasets are commonly used for generalization tests and were also used in most models evaluated here. In difference to most other evaluations that only report PCP, more metrics are used, and the PCP score is calculated as described above, and not by first averaging over the persons and then merging them together like usually, because this, on the one hand, results in a different weighting of errors since some persons occur more often than others, and on the other hand, allows a better comparison with other datasets.
In Table 3 it can be seen that the learnable approaches show similar performance on the Shelf dataset compared to the algorithmic ones. They are notably faster than most others, which mainly is caused by faster 2D pose estimators, but on the other hand, tend to predict more invalid persons. Note that in the dataset itself not all occurring persons are labeled, and therefore the percentage can not be zero, but the algorithmic models can detect all labeled persons while predicting a lower total number of persons. MvP shows an especially poor performance in this regard, even with the synthetically finetuned model, making it basically unusable for real-world applications (the preprocessed data and tools for Shelf & Campus were published so it could be trained here). While VoxelPose profits from the synthetic training, the difference for Faster-VoxelPose is only marginal or even slightly worse in some metrics. The camera setup of Shelf is quite similar to the Panoptic dataset, which is a reason the directly transferred models show a decent performance. In all cases, the models were evaluated end-to-end, instead of using the provided external heatmaps created by unpublished tools, since the generalization of the complete model is the interesting part here.
@@ 6 Using Depth Information
Figure 3: Example of depth masking in VoxelKeypointFusion. In the left image, the keypoint heatmap beam projection of four persons from five camera views into a voxelized room is shown, with one color for each joint type. This is the default input for the peak proposal calculation. The center shows the voxelized depth images, in which three persons (the fourth is walking through the entrance) and the room’s wall are clearly visible. On the right side, the projection was masked with the depth voxels. The three persons in the center of the room are now clearly visible in the projection space as well, and no peaks can be proposed between the persons anymore. Parts of the sphere’s wall are still left, but since the keypoint projections do not overlap there, no peak proposals are generated from those.
@@ 6.2 VoxelKeypointFusion
For VoxelKeypointFusion a simple depth-based masking concept was implemented. After generating a point cloud from the depth images, all points that fall into one voxel are counted, and if there are more than a small threshold (to drop noisy points), this voxel is considered filled. The projection space is then masked with the filled depth voxels, to drop all projection voxels that do not match to a real object. This especially helps to remove any voxel-peaks that are generated from 2D detections that do not show the same person but whose heatmap projections still overlap somewhere in 3D space. An example of the masking can be seen in Figure 3. The depth views do not have to be aligned to the color views, but can be recorded from any direction and also come from a different number of sensors.
@@ 8 Ablation studies
One limitation of VoxelKeypointFusion is, that since there is no neural refinement, the location accuracy is highly dependent on the voxel-resolution. As can be seen in Table 7, a lower voxel-resolution leads to better results, even though it does not directly scale (but $25mm$ is even better than the result with $50mm$ in Table 1). The voxel size also has a significant impact on the inference speed. As expected, with more cameras the results get better, but the inference time increases as well. With fewer cameras the performance drops, and the model predicts more invalid persons because overlaps where no persons stand are not pruned that much by different viewpoints that would see an empty space there. The number of persons, especially if there are many occlusions, has a notable impact on the inference speed as well. The total 3D time for Shelf is around $48ms$, for Campus around $56ms$, and for Panoptic, with fewer voxels but more total and occluded keypoints than Campus, around $100ms$.
@@ 10 Conclusion
VoxelKeypointFusion additionally shows that all the necessary person information is already present in the 2D heatmaps, which currently is ignored by the learnable voxel-based approaches. One limitation though is that since there is no neural refinement, the location accuracy is highly dependent on the voxel-resolution.